A food safety guideline is that the mercury in fish should be below 1 part per million​ (ppm). Listed below are the amounts of mercury​ (ppm) found in tuna sushi sampled at different stores in a major city. Construct a 99​% confidence interval estimate of the mean amount of mercury in the population. Does it appear that there is too much mercury in tuna​ sushi?
0.59  0.68  0.10  0.98  1.37  0.55  0.97
What is the confidence interval estimate of the population mean mu​?
_ppm < mu < _ ppm
​(Round to three decimal places as​ needed.)

Answer by Boreal(15235)

The mean is 5.24/7=0.749
sd=0.404
CI percentage not given. Will use 95%, although with mercury, one might want a higher confidence.
t 0.975,df=6 is 2.447
half-interval is t*s/sqrt(n)=2.447*0.404/sqrt(7)=0.374
95%CI is (0.375, 1.123) units ppm